If the second die always shows an odd no. find the conditional probability of getting a sum as 7,
if a pair of dice is to be known.

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
As second die always shows odd number 1,3,5 are possible
so number of possible ways=6*3=18 ways
also given sum should be 7 so possible ways are
(1,6) ,(3,4) ,(5,2) so there are 3 ways
So the conditional probability of getting a sum as 7,if a pair of dice is to be thrown
is 3/18=1/6